OpenStudy (rock_mit182):

Hi guys, I need help ! X and Y react according with the following equation: X (ac) + 2W (ac) --► Z ↓ + T (ac) taking 50 ml of a solution of X (0,1 M); and mixed up with 10 ml solution of 0,2 M of W. in accordance to the last mention. how much solution is left ?

OpenStudy (rock_mit182):

A. 0,003 moles de W B. 0,005 moles de X C. 0,001 moles de W D. 0,004 moles de X beyond the answer I need to clarify the approach , please someone help me out!

OpenStudy (cydney_morgan):

Hey its you!

OpenStudy (cydney_morgan):

Give me a second.. :)

OpenStudy (rock_mit182):

ok ;)

OpenStudy (cydney_morgan):

Sorry, I have not learned this yet, I'm trying to figure this out.... Give me a couple min... going into *insta* learn mode..

OpenStudy (rock_mit182):

well I'm good in math. but it seems to me that chemistry concepts allude my logic, hope you can help me out so take y

OpenStudy (rock_mit182):

your time*

OpenStudy (cydney_morgan):

Oh, I'm in the intensive math class but the honors science class... lol! I'm having issues sorry, I'll keep trying..

OpenStudy (rock_mit182):

impressive, you're the right one science girl ;)

OpenStudy (rock_mit182):

anyway i have till friday to come up with something so I'll give it a try also

OpenStudy (cydney_morgan):

Omg, I'm so sorry, my computer completely crashed.

OpenStudy (rock_mit182):

don't worry I'm going to read about the topic and then I'll figure it out, ;)

OpenStudy (cydney_morgan):

Thank you for not being mad!! I feel so bad I couldnt help. :'C

OpenStudy (rock_mit182):

I'm impressed with your commitment, that shows passion for what you do. when you need help in math, just tell me ..;)

OpenStudy (cydney_morgan):

Thank you so much, I'm taking a math test right now! I'll tell you if I need help. :)

OpenStudy (rock_mit182):

ok.

OpenStudy (aaronq):

Find moles using the molarity and volume; \(M=\dfrac{n_{solute}}{L_{solution}}\) subtract one from the other taking into account the stoichiometric coefficients.

OpenStudy (rock_mit182):

\[[(0,01L)*(0,2M)= 2 solute ] - [(0,05 L) * (0,1M) = x solute] \]

OpenStudy (rock_mit182):

is that what you mean ?

OpenStudy (aaronq):

yeah, the units of the equation are moles (because molarity is "mol/L", so L cancels out)

OpenStudy (rock_mit182):

thanks a lot !, are you proficient in general chemistry ?

OpenStudy (aaronq):

no problem. yes i am

OpenStudy (rock_mit182):

good to know, the math is no so hard but I struggle with the concepts

OpenStudy (aaronq):

yeah, the math is arithmetic.The concepts take a little logic, it helps to picture what is going on in real life (as if you were in the lab).

OpenStudy (rock_mit182):

I'll take the advice, appreciate that